OpenStudy (anonymous):

Find an equation of the perpendicular bisector of the line segment with endpoints (1,4) & (3,10). (Can be done either distance formula or endpoints ( mid-point equation.)

OpenStudy (anonymous):

is this right? y-4=3(x-1)

OpenStudy (anonymous):

Bisector = intersects in middle. Point of intersection = (2, 7)

OpenStudy (anonymous):

That is the equation for the line between 1,4 and 3,10. Right. Now, you need the equation of the line perpendicular to it, and passes through 2,7.

OpenStudy (anonymous):

but how can I find that?

OpenStudy (anonymous):

Slope of perpendicular lines multiply to -1. So, the perpendicular line's slop will be -1/3.

OpenStudy (anonymous):

oh so u mean m=3 which turns out m=-1/3?

OpenStudy (anonymous):

m=slope..

OpenStudy (anonymous):

correct. m1 = 3. m2 = -1/3.

OpenStudy (anonymous):

gotch yaa.. so the equation for the perpendicular bisector is: y-4=-1/3(x-1)

OpenStudy (anonymous):

No. This line goes through 2,7.